A rectangular coil with sides 0.150 m by 0.210 m has 505 turns of wire. It is rotated about its long axis in a uniform magnetPlease include explanation
Answer #1

Umax= NBAW- NBA (27f) Vmax - frequency, f ) NBA( 2A) - f= 115 505 x 0.55× 0,15 ×0,21 X 2 x 3,14 f = 2.093 H2 2.09 Hz ſ